OpenStudy (anonymous):

Arnold went shopping at The Link with $80 to spend. He spent $12 on a belt and $20 for a shirt. wat is an inequality to represent how much he can spend on pants, x.

OpenStudy (anonymous):

i need a awnser

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Do you have a guess?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok well he had $80, then he spent a total of $32 ($12 of a belt and $20 on a shirt). If he wanted to buy a pair of pants with the remaining money (x) it would have to be equal to or less than what he has left, right? Is this making sense?

OpenStudy (anonymous):

no i dont get it do you think we can get some help from somebody you know.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

Yeah...I'm not sure how I could explain this differently. I'll give a shout-out to some people. @phi @AnswerMyQuestions @Whitemonsterbunny17 @iGreen @Cosmichaotic Do you guys think you could help him with this?

OpenStudy (anonymous):

So $80 minus what he spent on the belt and shirt leaves him with how much?

OpenStudy (anonymous):

$48 right? So if x is the cost of pants. Then \[x \le $48\]

OpenStudy (anonymous):

That would be how much he has to spend on pants. Considering there is no tax.

OpenStudy (igreen):

\(80 - 20 - 12 \le x\) Where x = Pants

OpenStudy (igreen):

Simplify: \(60 - 12 \le x\) \(48 \le x\)

OpenStudy (igreen):

@rashional Can you give me a medal? :l

OpenStudy (anonymous):

I'll give you one for explaining it properly. Thanks!

OpenStudy (igreen):

Oh, ty. @Cosmichaotic

OpenStudy (anonymous):

wats the inequality

OpenStudy (igreen):

Both me and @Cosmichaotic posted it.

OpenStudy (anonymous):

so its 60 -12 and the other sign

OpenStudy (igreen):

No. \(x \le 48\)

OpenStudy (anonymous):

thats it

OpenStudy (igreen):

Yes